OpenStudy (anonymous):

Derek distributed twenty-two pieces of apple to his friends. If each apple slice was one-eighth of an apple, how many apples did he distribute to his friends? three and one-fourth two and three-fourths three and one-eighth two and three-eighths

OpenStudy (anonymous):

\[1/8 \times 22\] = \[2\frac{3}{4}\]

OpenStudy (diyadiya):

Let the number of apples be x 1 piece = 1/8 of one apple =1/8*x 22 piece= 1/8 * x *22 =22x/8 \[=\frac{11}{4} = 2 \frac{3}{4}\]

OpenStudy (anonymous):

thankz